OpenStudy (anonymous):

quadratic equations : solve: (x+1)^2 = 16?

OpenStudy (anonymous):

\[x + 1 = \pm 4\]

OpenStudy (anonymous):

x + 1 = 4 x = 3 x + 1 = -4 x = -5

OpenStudy (anonymous):

Do you want to know the steps he followed?